OpenStudy (anonymous):

given that g(x)=2-3x and h(x)=mx+n, where m and n are constant. find a) g^-1 (x) b) g^2 (x) c) the value of m and of n if gh(x) = -x-1

OpenStudy (mani_jha):

The first one is obtained by getting the reciprocal of g(x).(1/(2-3x) Square g(x) to get the second one. Put x = mx+n in g(x) g(h(x))=2-3(mx+n)=-x-1 Maybe I interpreted the question incorrectly, but you cannot solve both for m and n. Is there any other information given?

OpenStudy (anonymous):

no~